OpenStudy (anonymous):

Multiply and simplify: k^8 • k^–8 A. 0 B. 1 C. k D. 8k

ganeshie8 (ganeshie8):

\(\large a^m . a^n = a^{m+n}\)

OpenStudy (anonymous):

What do you think it is raster given what ganeshie said?

OpenStudy (anonymous):

k^8 • k^–8 = k^x • k^–x ; )

OpenStudy (anonymous):